OpenStudy (anonymous):

Solve for x: 5x + 2 = 4x - 9. 7 -7 11 -11

OpenStudy (anonymous):

5x + 2 = 4x - 9. 5x-4x =-9-2 x= ?

OpenStudy (anonymous):

sorry x=-11